OpenStudy (anonymous):

Prove that cos2 2x – cos2 6x = sin 4x sin 8x

OpenStudy (tyteen4a03):

\(cos^2(2x) - cos^2(6x) = sin(4x)sin(8x)\) Hint: There are two identities you need: 1. Power reduction: \(\large{cos^2(\theta) = \frac{1+cos(2\theta)}{2}}\) 2. Product-to-sum identity: \(\large{sin(x)sin(y) = \frac{1}{2} (cos(x - y) - cos(x + y))}\)